OpenStudy (anonymous):

domain and range of square root of x+6

jimthompson5910 (jim_thompson5910):

Is it \[\Large \sqrt{x+6}\] or is it \[\Large \sqrt{x}+6\] ?? Also, how far did you get with this problem?

OpenStudy (anonymous):

1st one. I figured x=/= -6, but i'm not sure if the domain and range would be unions or not.

OpenStudy (samsan9):

i got this question and i believe the domain is (-6,infinity) and (0,infinity) i could be wrong though...

jimthompson5910 (jim_thompson5910):

well let's see what happens when x = -6 \[\Large \sqrt{x+6}\] \[\Large \sqrt{-6+6}\] \[\Large \sqrt{0}\] \[\Large 0\] So x = -6 is certainly possible and allowed in the domain.

OpenStudy (anonymous):

oh, because it has to be either 0 or a positive number?

jimthompson5910 (jim_thompson5910):

It turns out that the radicand (basically the stuff under the square root) cannot be negative. So... \[\Large x + 6 \ge 0\] \[\Large x + 6{\color{red}-6} \ge 0{\color{red}-6}\] \[\Large x + 0 \ge -6\] \[\Large x \ge -6\]

jimthompson5910 (jim_thompson5910):

the radicand, yes, it must be 0 or positive

OpenStudy (anonymous):

so, the range is x is greater than or equal to -6? How would I figure out the range?

jimthompson5910 (jim_thompson5910):

What's the smallest value in the domain?

OpenStudy (anonymous):

-6?

jimthompson5910 (jim_thompson5910):

plug that into the function to get ???

OpenStudy (anonymous):

ohh! square root of 0?

jimthompson5910 (jim_thompson5910):

which is ???

OpenStudy (anonymous):

0

jimthompson5910 (jim_thompson5910):

So that's the smallest value in the range. Smallest value in domain ----> smallest value in the range

jimthompson5910 (jim_thompson5910):

This only works because the function sqrt(x+6) is always increasing.

OpenStudy (anonymous):

I get it, thank you :P

jimthompson5910 (jim_thompson5910):

you're welcome